a^2=-27-12a PLEASE HELP!
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

it's a "quadratic", to solve move everything to one side:
a^2=-27-12a
a^2+12a=-27
a^2+12a+27=0
factor the left:
(a+3)(a+9) = 0
a = {-3, -9}

Answer by stanbon(75887)

a^2=-27-12a
---------------
Rearrange:
a^2+12a+27 = 0
Factor:
a^2+3a+9a+27 = 0
a(a+3)+9(a+3) = 0
-----
Factor:
(a+3)(a+9) = 0
a = -3 or a = -9
======================
Cheers,
Stan H.
============